# How to Securely Shuffle? A survey about Secure Shufflers for privacy-preserving computations

### v1

@@ 4.2.3. Noticeable solution: Tor as a Shuffler (TORaaS)
While this solution is attractive due to its simplicity, practitioners should carefully consider the possibility of de-anonymization attacks (Arp et al., 2015). Even though there exists mitigation techniques, Tor is still particularly sensitive to such attacks as it is a free-route mix network.
@@ 4.3. Dining Cryptographers networks (DC nets)
These solutions take inspiration from the Dining Cryptographers problem (Chaum, 1988; Golle and Juels, 2004). This anonymity problem is named after a scenario where three cryptographers are dining together and learn that their meal has been paid for, but they want to know whether one of them paid or if it was the NSA; without revealing who paid if it was one of them.
The solution proposed by Chaum (1988) is to exchange complementary binary masks between the participants (as represented in Fig. 3). Then, each participant publishes a masked value. Once aggregated, the masked values reveal the anonymous bit of information. This masking trick was reused as a foundation for several anonymous protocols referred to as “DC nets”. Most anonymization systems based on DC nets complement DC nets with other cryptographic solutions to combine scalability and security. Recently, Shirali et al. (Shirali et al., 2022) wrote a survey focused on DC nets.
Figure 3. Dining Cryptographers problem
[…]
[Schema of the DC problem]Schema of the DC problem: three secret owners exchange additive masks with each other and then publish their mask values to reveal the final result.
We identified four major DC nets: Dissent (Wolinsky et al., 2012), Riposte (Corrigan-Gibbs et al., 2015), Spectrum (Newman et al., 2022), and Sabre (Vadapalli et al., 2022).
@@ 4.3.1. Scalability
Until 2012 (Wolinsky et al., 2012), DC nets had been considered a purely theoretical construction compared to the highly scalable mix networks. Dissent (Wolinsky et al., 2012) marked a turning point because practical applications could be imagined, but scalability was still very limited. Recent solutions, such as Riposte (Corrigan-Gibbs et al., 2015) and Spectrum (Newman et al., 2022), are the first scalable DC nets. Riposte was the first anonymous broadcaster able to shuffle 1M messages within a few hours. It is not as scalable as mix networks but requires fewer servers (i.e., 2 or 3).
@@ 4.3.2. Security
Each DC net has a rather unique architecture, so they do not share many common points, except their foundation: the DC problem. For simplicity, we only highlight the properties of the state-of-the-art method: Spectrum (Newman et al., 2022). It is most optimized for two servers and is anonymous against one malicious server. There are no known de-anonymization attacks specific to Spectrum (or DC nets in general). Spectrum has correctness and is resistant to sender disruption only.
@@ 4.4. Verifiable shufflers
For more details about verifiable shufflers, we recommend the survey paper of (Haines and Müller, 2020).
Figure 4. Classic verifiable shufflers
[…]
[Schema of a verifiable shuffler]Schema of a verifiable shuffler: three data owners send their encrypted message to a chain of servers, the servers successively shuffles the ciphertexts, prove the shuffle and verify each other proofs. Finally, they transmit the final output to an analyzer.
We identify four major verifiable shufflers in the literature: Classic verifiable shufflers (Bayer and Groth, 2012), Riffle (Kwon et al., 2016), Atom (Kwon et al., 2017), lattice-based verifiable shuffle (Aranha et al., 2023).
We use Bayer and Groth (2012) as reference for classic verifiable shuffling, but some recent works (Hébant et al., 2020; Fauzi et al., 2017) have described various optimizations such as homomorphic proofs (Hébant et al., 2020) (making the verification cost constant in the number of nodes in the network).
Recently, Balle et al. (2023) used verifiable shufflers to build an “alternating shuffler.” This work split the messages to shuffle into several subsets and shuffle them independently. An alternating shuffler outputs a permutation that is imperfect, but good enough to provide DP guarantees.
@@ 4.4.1. Scalability
The classic verifiable shufflers (Bayer and Groth, 2012) have some scalability bottlenecks (i.e., cannot scale above tens of thousands of messages), but it does not prevent concrete use cases. However, the most recent adaptations of verifiable shufflers to anonymous communications, namely Riffle (Kwon et al., 2016) and Atom (Kwon et al., 2017), do not have this scalability issue. For example, Atom can shuffle a million message in less than an hour.
@@ 4.5. MPC-based shufflers
Figure 5. MPC-based secure shufflers ($\pi$: an MPC protocol)
[…]
[Schema of an MPC-based shuffler]Schema of an MPC-based shuffler: three data owner send one share of their secret to three different servers. The servers then perform an MPC shuffle program and transmit the output to an analyzer.
We identified six major MPC-based shufflers in the literature: AsynchroMix (Lu et al., 2019b), PowerMix (Lu et al., 2019b), Blinder (Abraham et al., 2020), Clarion (Eskandarian and Boneh, 2022), RPM (Lu and Kate, 2023), and Ruffle (A et al., 2023).
Some of these shufflers (e.g., Clarion) are built upon generic “secret-shared shuffle protocols.” This generic building block could be further optimized thanks to recent advances (Chase et al., 2020; Gao et al., 2024; Song et al., 2024).
@@ 4.5.2. Security
All these techniques (except Clarion) share similar security properties: anonymous if there is a majority of honest servers, correctness, and sender disruption resistance. Blinder even has server disruption resistance. On the other hand, Clarion only has sender disruption resistance (i.e., no correctness and no server disruption resistance), but it is anonymous against $M-1$ malicious servers. No de-anonymization attack specific to these solutions has been yet described.
@@ 4.6. TEE-based shufflers
Bittau et al. (Bittau et al., 2017) introduced the first TEE-based secure shuffler. The only step to conceive a TEE-based secure shuffler is to provide a shuffle code executable on the chosen architecture (e.g., Intel SGX (Intel, 2015)). Shuffling requires simple operations, but the implementation must be oblivious: the server owner should not be able to infer information about the data based on the memory access patterns. In other words, TEE-based shufflers behave as a proxy in a secure environment that guarantees data confidentiality. We identified four TEE-based secure shufflers: Prochlo (Bittau et al., 2017), ORShuffle (Sasy et al., 2022), the Butterfly Oblivious Shuffle (Gu et al., 2023), and DBucket (Ngai et al., 2024).
Figure 6. TEE-based secure shufflers
[…]
[Schema of a TEE-based shuffler]Schema of a TEE-based shuffler: three data owner send their encrypted message to the secure enclave of a server, a shuffle is performed in the secure enclave and the result is transmitted to an analyzer.
@@ 4.6.2. Security
Moreover, flawed TEE implementations can be sensitive to side-channel attacks. For example, Sasy et al. (Sasy et al., 2022) highlighted that Prochlo is sensitive to some recent attacks.
@@ 4.7. Other shufflers
Finally, we identified three other secure shufflers that do not fit in any of the previous categories: rumor-based shuffler (Fanti et al., 2017), NIDAR (Non-Interactive Differentially Anonymous Router) (Bünz et al., 2021), and secure-aggregation-based shuffler (Bell et al., 2020).
@@ 4.8. Scalability with the message size
Even if mix networks use public-key encryption [^1] Symmectric encryption has a linear computation cost in the message size while public-key encryption induces a cubic computation cost in the message size., they all have a linear dependency on message size. In practice, public-key encryption in mix networks is only used to agree on a symmetric key used by the communicating parties for lightweight encryption and decryption.
The computation costs of proof and verification protocols usually correspond to the number of modular exponentiations needed. Since modular exponentiations have a cubic cost, it induces a $O(L^{3})$ dependency in classic verifiable shufflers (Bayer and Groth, 2012) and Atom (Kwon et al., 2017).
All the DC-nets have a linear dependency on the message size. For Riposte, Spectrum, and Sabre, an additive term (linear in the message size) appears because of a cryptographic primitive they use (i.e., Function Secret Sharing (Gilboa and Ishai, 2014)). All the TEE-based and MPC-based shufflers have a linear dependency on the message size. Among the other shufflers, only secure-aggregation-based shufflers have a non-linear dependency on message size.
@@ 5.1. Private statistics
Following the first anonymous communication systems, several works (Brickell and Shmatikov, 2006; Durak et al., 2023; Günther et al., 2020; Hohenberger et al., 2014) have built systems based on message anonymization to compute private aggregate statistics (e.g., mean). As detailed in Section 3, the sender anonymization layer used in these work can be seen as a shuffler. This equivalence between message anonymization and shuffling is common in Differential Privacy where papers alternatively use “anonymous messages” (Ghazi et al., 2021, 2020c) and “shuffled messages” (Girgis et al., 2020; Erlingsson et al., 2020).
On the one hand, some systems (Brickell and Shmatikov, 2006; Hohenberger et al., 2014) relied on shuffler to anonymize the message sender. For example, Brickell et al. (Brickell and Shmatikov, 2006) built a privacy-preserving data collection system based on such message anonymization. In their system, each data owner sends their private data to the shuffler and the shuffler outputs a set of “anonymous” messages. This approach provides a weak privacy preservation as it does not mitigate any privacy leakage caused by the message content.
On the other hand, other works (Durak et al., 2023; Günther et al., 2020) built inspired by Ishai et al. (Ishai et al., 2006): they require each data owner to generate secret shares and to send these shares to the shuffler. As explained in Section 2.4, the secret sharing will prevent the secret reconstruction. Thus, these works provided stronger privacy guarantees than (Brickell and Shmatikov, 2006; Hohenberger et al., 2014) because they mitigate any privacy leakage from individual messages; only the aggregated result is revealed.
@@ 5.2. Differentially Private computations
While private statistics systems such as (Durak et al., 2023; Günther et al., 2020) already provide strong privacy guarantees, works in DP proposes an even stronger privacy amplification mitigating privacy leakage both from the shuffled messages and from the aggregated output.
This DP framework based on shuffler is called “the shuffle DP model” (Ghazi et al., 2020b) and consists of two randomized algorithms (see Fig. 7):
@@
Figure 7. Shuffle Differential Privacy: (1) The data owners split their private value into several random messages using the randomizer $\mathcal{R}$, (2) they transmit these messages to $\mathcal{S}$, (3) the shuffler outputs a random permutation of the messages and (4) the analyser can perform any relevant operation on these messages.
[…]
[Schema of the Shuffle DP framework]Schema of the Shuffle DP framework: three private data samples are each split into 3 messages by a local randomizer. The messages are transmitted to a shuffler that shuffles them and transmit the shuffle output to an analyzer.
@@
Several papers (Balle et al., 2019b; Chen et al., 2024; Erlingsson et al., 2019; Feldman et al., 2023) showed that the shuffling induces a privacy amplification because it requires less noise than local DP to offer comparable privacy guarantees. The privacy-accuracy trade-off of Shuffle DP approaches the trade-off of Central DP without requiring a trusted party. This good trade-off without a trusted party assumption often motivates the use of shuffle DP in privacy-preserving computations. For an extensive overview on the benefits of Shuffle DP compared to local and central DP, we refer to Cheu (Cheu, 2021).
While shuffle DP is usable for any statistical computation, it provides a privacy-utility close to central DP mostly for sum operations. The first shuffle DP papers focused on private summation (Balle et al., 2019a, 2020a; Ghazi et al., 2020a), but recent works described more complex statistical operations: private quantiles (Aamand et al., 2025), private uniformity testing (Balcer et al., 2021; Cheu, 2021), private histograms (Balcer and Cheu, 2020; Bell et al., 2022a; Wang et al., 2025) and other related computations (Ghazi et al., 2021; Luo et al., 2022; Humphries et al., 2022).
@@ 5.3. Federated Learning and Secure Aggregation
Federated Learning (FL) (Kairouz et al., 2021) is a decentralized ML paradigm enabling to collaboratively train an ML model without exchanging their (private) training data. While there exists several FL variants, the most popular setup is “horizontal FL:” each data owner owns a small dataset over the same set of features.
@@
Figure 8. High-level Federated Learning architecture
[…]
[Schema of an FL architecture]Schema of an FL architecture: multiple data owners each trains locally an ML model and then communicate with a group of coordination servers outputting a global ML model
Such aggregation-based FL protocols enable to train a wide variety of ML models: any model which training relies on gradient descent-based (Konečný et al., 2016) (e.g., neural networks or linear regression), k-means protocols (Zhang et al., 2022; Li et al., 2023; Holzer et al., 2023; Kumar et al., 2020), Singular Vector Decomposition (Chai et al., 2021), random forests (de Souza et al., 2020; Liu et al., 2020), and Gradient Boosted Decision Trees (Tian et al., 2020; Cheng et al., 2021).
Even though FL does not require private data exchanges, the ML models (or gradients) exchanged during these protocols can leak private information (Nasr et al., 2019; Dartel et al., 2025). To enhance privacy, many works (Kairouz et al., 2021) have relied on secure aggregation to obtain the sum of the local models without revealing the individual models.
@@
Secure aggregation is a protocol taking as input $n$ private vectors $v_{i}$ and outputting $\sum_{i=1}^{n}v_{i}$ without revealing the input vectors. Several solutions have been proposed to implement secure aggregation including additive masking (Bonawitz et al., 2017; Bell et al., 2020; Sabater et al., 2021; Bell et al., 2022b), Homomorphic Encryption (Shi et al., 2011), Trusted-Execution Environment (TEE) (Lie and Maniatis, 2017), and shuffling-based (Ishai et al., 2006). Mansouri et al. (2023) surveyed secure aggregation protocols.
We can implement shuffling-based secure aggregation in two manners. On the one hand, we can perform an “exact” secure aggregation using the protocol proposed by Ishai et al. (2006) (see in Section 2.4). On the other hand, we can perform a “differentially private” secure aggregation using the DP summation mentioned in Section 5.2. Several FL papers have recently used this shuffle-DP aggregation (Girgis et al., 2021; Liu et al., 2021; Xu et al., 2024).
@@
There is a growing interest in gradient sparsification to reduce communications in FL (Stich et al., 2018). As for classic secure aggregation, the sparse secure aggregation can replace the non-private aggregation step to enhance privacy. Luo et al. (Luo et al., 2022) presented a shuffling-based sparse vector aggregation.

### v2

@@ 4.2.3. Noticeable solution: Tor as a Shuffler (TORaaS)
While this solution is attractive due to its simplicity, practitioners should carefully consider the possibility of de-anonymization attacks (arp_torben_2015). Even though there exists mitigation techniques, Tor is still particularly sensitive to such attacks as it is a free-route mix network.
@@ 4.3. Dining Cryptographers networks (DC nets)
These solutions take inspiration from the Dining Cryptographers problem (chaum_dining_1988; golle_dining_2004). This anonymity problem is named after a scenario where three cryptographers are dining together and learn that their meal has been paid for, but they want to know whether one of them paid or if it was the NSA; without revealing who paid if it was one of them.
The solution proposed by chaum_dining_1988 is to exchange complementary binary masks between the participants (as represented in Fig. 3). Then, each participant publishes a masked value. Once aggregated, the masked values reveal the anonymous bit of information. This masking trick was reused as a foundation for several anonymous protocols referred to as “DC nets”. Most anonymization systems based on DC nets complement DC nets with other cryptographic solutions to combine scalability and security. Recently, Shirali et al. (shirali_survey_2022) wrote a survey focused on DC nets.
Figure 3. Dining Cryptographers problem: Schema of the DC problem
We identified four major DC nets: Dissent (wolinsky_dissent_2012), Riposte (corrigan-gibbs_riposte_2015), Spectrum (newman_spectrum_2022), and Sabre (vadapalli_sabre_2022).
@@ 4.3.1. Scalability
Until 2012 (wolinsky_dissent_2012), DC nets had been considered a purely theoretical construction compared to the highly scalable mix networks. Dissent (wolinsky_dissent_2012) marked a turning point because practical applications could be imagined, but scalability was still very limited. Recent solutions, such as Riposte (corrigan-gibbs_riposte_2015) and Spectrum (newman_spectrum_2022), are the first scalable DC nets. Riposte was the first anonymous broadcaster able to shuffle 1M messages within a few hours. It is not as scalable as mix networks but requires fewer servers (i.e., 2 or 3).
@@ 4.3.2. Security
Each DC net has a rather unique architecture, so they do not share many common points, except their foundation: the DC problem. For simplicity, we only highlight the properties of the state-of-the-art method: Spectrum (newman_spectrum_2022). It is most optimized for two servers and is anonymous against one malicious server. There are no known de-anonymization attacks specific to Spectrum (or DC nets in general). Spectrum has correctness and is resistant to sender disruption only.
@@ 4.4. Verifiable shufflers
For more details about verifiable shufflers, we recommend the survey paper of (haines_sok_2020).
Figure 4. Classic verifiable shufflers: Schema of a verifiable shuffler
We identify four major verifiable shufflers in the literature: Classic verifiable shufflers (bayer_efficient_2012), Riffle (kwon_riffle_2016), Atom (kwon_atom_2017), lattice-based verifiable shuffle (aranha_verifiable_2023).
We use bayer_efficient_2012 as reference for classic verifiable shuffling, but some recent works (hebant_linearly-homomorphic_2020; fauzi_efficient_2017) have described various optimizations such as homomorphic proofs (hebant_linearly-homomorphic_2020) (making the verification cost constant in the number of nodes in the network).
Recently, balle_amplification_2023 used verifiable shufflers to build an “alternating shuffler.” Their protocol splits the messages to shuffle into several subsets and shuffle them independently. An alternating shuffler outputs a permutation that is imperfect, but good enough to provide DP guarantees.
@@ 4.4.1. Scalability
The classic verifiable shufflers (bayer_efficient_2012) have some scalability bottlenecks (i.e., cannot scale above tens of thousands of messages), but it does not prevent concrete use cases. However, the most recent adaptations of verifiable shufflers to anonymous communications, namely Riffle (kwon_riffle_2016) and Atom (kwon_atom_2017), do not have this scalability issue. For example, Atom can shuffle a million message in less than an hour.
@@ 4.5. MPC-based shufflers
Figure 5. MPC-based secure shufflers ($\pi$: an MPC protocol): Schema of an MPC-based shuffler
We identified six major MPC-based shufflers in the literature: AsynchroMix (lu_honeybadgermpc_2019), PowerMix (lu_honeybadgermpc_2019), Blinder (abraham_blinder_2020), Clarion (eskandarian_clarion_2022), RPM (lu_rpm_2023), and Ruffle (a_ruffle_2023).
Some of these shufflers (e.g., Clarion) are built upon generic “secret-shared shuffle protocols.” This generic building block could be further optimized thanks to recent advances (chase_secret-shared_2020; gao_multiparty_2024; song_secret-shared_2024).
@@ 4.5.2. Security
All these techniques (except Clarion) share similar security properties: anonymous if there is a majority of honest servers, correctness, and sender disruption resistance. Blinder even has server disruption resistance. On the other hand, Clarion only has sender disruption resistance (i.e., no correctness and no server disruption resistance), but it is anonymous against M1M-1 malicious servers. No de-anonymization attack specific to these solutions has been yet described.
@@ 4.6. TEE-based shufflers
Bittau et al. (bittau_prochlo_2017) introduced the first TEE-based secure shuffler. The only step to conceive a TEE-based secure shuffler is to provide a shuffle code executable on the chosen architecture (e.g., Intel SGX (intel_software_2015)). Shuffling requires simple operations, but the implementation must be oblivious: the server owner should not be able to infer information about the data based on the memory access patterns. In other words, TEE-based shufflers behave as a proxy in a secure environment that guarantees data confidentiality. We identified four TEE-based secure shufflers: Prochlo (bittau_prochlo_2017), ORShuffle (sasy_fast_2022), the Butterfly Oblivious Shuffle (gu_efficient_2023), and DBucket (ngai_distributed_2024).
Figure 6. TEE-based secure shufflers: Schema of a TEE-based shuffler
@@ 4.6.2. Security
Moreover, flawed TEE implementations can be sensitive to side-channel attacks. For example, Sasy et al. (sasy_fast_2022) highlighted that Prochlo is sensitive to some recent attacks.
@@ 4.7. Other shufflers
Finally, we identified three other secure shufflers that do not fit in any of the previous categories: rumor-based shuffler (fanti_hiding_2017), NIDAR (Non-Interactive Differentially Anonymous Router) (bunz_non-interactive_2021), and secure-aggregation-based shuffler (bell_secure_2020).
@@ 4.8. Scalability with the message size
The computation costs of proof and verification protocols usually correspond to the number of modular exponentiations needed. Since modular exponentiations have a cubic cost, it induces a $O(L^{3})$ dependency in classic verifiable shufflers (bayer_efficient_2012) and Atom (kwon_atom_2017).
All the DC-nets have a linear dependency on the message size. For Riposte, Spectrum, and Sabre, an additive term (linear in the message size) appears because of a cryptographic primitive they use (i.e., Function Secret Sharing (gilboa_distributed_2014)). All the TEE-based and MPC-based shufflers have a linear dependency on the message size. Among the other shufflers, only secure-aggregation-based shufflers have a non-linear dependency on message size.
@@ 5.1. Private statistics
Following the first anonymous communication systems, several works (brickell_efficient_2006; durak_precio_2023; gunther_pem_2020; hohenberger_anonize_2014) have built systems based on message anonymization to compute private aggregate statistics (e.g., mean). As detailed in Section 3, the sender anonymization layer used in these work can be seen as a shuffler. This equivalence between message anonymization and shuffling is common in Differential Privacy where papers alternatively use “anonymous messages” (ghazi_power_2021; ghazi_private_2020) and “shuffled messages” (girgis_shuffled_2020; erlingsson_encode_2020).
On the one hand, some systems (brickell_efficient_2006; hohenberger_anonize_2014) relied on shuffler to anonymize the message sender. For example, Brickell et al. (brickell_efficient_2006) built a privacy-preserving data collection system based on such message anonymization. In their system, each data owner sends their private data to the shuffler and the shuffler outputs a set of “anonymous” messages. This approach provides a weak privacy preservation as it does not mitigate any privacy leakage caused by the message content.
On the other hand, other works (durak_precio_2023; gunther_pem_2020) built inspired by Ishai et al. (ishai_cryptography_2006): they require each data owner to generate secret shares and to send these shares to the shuffler. As explained in Section 2.4, the secret sharing will prevent the secret reconstruction. Thus, these works provided stronger privacy guarantees than (brickell_efficient_2006; hohenberger_anonize_2014) because they mitigate any privacy leakage from individual messages; only the aggregated result is revealed.
@@ 5.2. Differentially Private computations
While private statistics systems such as (durak_precio_2023; gunther_pem_2020) already provide strong privacy guarantees, works in DP proposes an even stronger privacy amplification mitigating privacy leakage both from the shuffled messages and from the aggregated output.
This DP framework based on shuffler is called “the shuffle DP model” (ghazi_private_2020-1) and consists of two randomized algorithms (see Fig. 7):
@@
Figure 7. Shuffle Differential Privacy: (1) The data owners split their private value into several random messages using the randomizer $\mathcal{R}$, (2) they transmit these messages to $\mathcal{S}$, (3) the shuffler outputs a random permutation of the messages and (4) the analyser can perform any relevant operation on these messages.: Schema of the Shuffle DP framework
@@
Several papers (balle_privacy_2019; chen_generalized_2024; erlingsson_amplification_2019; feldman_stronger_2023) showed that the shuffling induces a privacy amplification because it requires less noise than local DP to offer comparable privacy guarantees. The privacy-accuracy trade-off of Shuffle DP approaches the trade-off of Central DP without requiring a trusted party. This good trade-off without a trusted party assumption often motivates the use of shuffle DP in privacy-preserving computations. For an extensive overview on the benefits of Shuffle DP compared to local and central DP, we refer to Cheu (cheu_differential_2021).
While shuffle DP is usable for any statistical computation, it provides a privacy-utility close to central DP mostly for sum operations. The first shuffle DP papers focused on private summation (balle_improved_2019; balle_private_2020; ghazi_pure_2020), but recent works described more complex statistical operations: private quantiles (aamand_lightweight_2025), private uniformity testing (balcer_connecting_2021; cheu_differential_2021), private histograms (balcer_separating_2020; bell_distributed_2022; wang_beyond_2025) and other related computations (ghazi_power_2021; luo_frequency_2022; humphries_selective_2022).
@@ 5.3. Federated Learning and Secure Aggregation
Federated Learning (FL) (kairouz_advances_2021) is a decentralized ML paradigm enabling to collaboratively train an ML model without exchanging their (private) training data. While there exists several FL variants, the most popular setup is “horizontal FL:” each data owner owns a small dataset over the same set of features.
@@
Figure 8. High-level Federated Learning architecture: Schema of an FL architecture
Such aggregation-based FL protocols enable to train a wide variety of ML models: any model which training relies on gradient descent-based (konecny_federated_2016) (e.g., neural networks or linear regression), k-means protocols (zhang_practical_2022; li_differentially_2023; holzer_dynamically_2023; kumar2020federated), Singular Vector Decomposition (chai_federated_2021), random forests (de_souza_dfedforest_2020; liu_federated_2020), and Gradient Boosted Decision Trees (tian_federboost_2020; cheng_secureboost_2021).
Even though FL does not require private data exchanges, the ML models (or gradients) exchanged during these protocols can leak private information (nasr_comprehensive_2019; dartel_evaluating_2025). To enhance privacy, many works (kairouz_advances_2021) have relied on secure aggregation to obtain the sum of the local models without revealing the individual models.
@@
Secure aggregation is a protocol taking as input $n$ private vectors $v_{i}$ and outputting $\sum_{i=1}^{n}v_{i}$ without revealing the input vectors. Several solutions have been proposed to implement secure aggregation including additive masking (bonawitz_practical_2017; bell_secure_2020; sabater_accurate_2021; bell_acorn_2022), Homomorphic Encryption (shi_privacy-preserving_2011), Trusted-Execution Environment (TEE) (lie_glimmers_2017), and shuffling-based (ishai_cryptography_2006). mansouri_sok_2023 surveyed secure aggregation protocols.
We can implement shuffling-based secure aggregation in two manners. On the one hand, we can perform an “exact” secure aggregation using the protocol proposed by ishai_cryptography_2006 (see in Section 2.4). On the other hand, we can perform a “differentially private” secure aggregation using the DP summation mentioned in Section 5.2. Several FL papers have recently used this shuffle-DP aggregation (girgis_shuffled_2021; liu_flame_2021; xu_camel_2024).
@@
There is a growing interest in gradient sparsification to reduce communications in FL (stich_sparsified_2018). As for classic secure aggregation, the sparse secure aggregation can replace the non-private aggregation step to enhance privacy. Luo et al. (luo_frequency_2022) presented a shuffling-based sparse vector aggregation.